# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

### v3

@@ Appendix A IMM filter theory
IMM-JHSE runs two ground plane EKFs in parallel, each with the state vector definition in Equation 13. One uses the static homography dynamics model in Equation 11, while the other uses the dynamic one in Equation 10. The IMM methodology allows the proposed method to switch between and mix state estimates from both filters, proportional to their likelihood [80]. This way, IMM-JHSE is robust to scenes with both static and dynamic camera movement. This section briefly describes the operation of IMM filters according to [77, 79, 78, 80], which is implemented with the FilterPy library [90].
@@
The mixing probabilities $\mu_{t|t-1}^{i|j}$ represent the probability of transitioning from model $j$ at time $t-1$ to model $i$ at time $t$. These are computed using the Markov transition probabilities $p_{ji}$:
@@ Appendix B Deriving the Jacobian matrix
[…]

### v4

@@ Appendix A IMM filter theory
IMM-JHSE runs two ground plane EKFs in parallel, each with the state vector definition in Equation 13. One uses the static homography dynamics model in Equation 11, while the other uses the dynamic one in Equation 10. The IMM methodology allows the proposed method to switch between and mix state estimates from both filters, proportional to their likelihood (Blom1988TheCoefficients). This way, IMM-JHSE is robust to scenes with both static and dynamic camera movement. This section briefly describes the operation of IMM filters according to (imm; imm_genovese; Blackman1999DesignSystems; Blom1988TheCoefficients), which is implemented with the FilterPy library (filterpy).
@@
The mixing probabilities $\mu_{t|t-1}^{i|j}$ represent the probability of transitioning from model $j$ at time t1t-1 to model $i$ at time $t$. These are computed using the Markov transition probabilities $p_{ji}$:
@@ Appendix B Deriving the Jacobian matrix
$\mathbf{b}=\begin{bmatrix}b_{1}\\ b_{2}\\ b_{3}\end{bmatrix}=\mathbf{H}^{G}_{t}\begin{bmatrix}x^{W}_{t}\\ y^{W}_{t}\\ 1\end{bmatrix}$,